OpenStudy (anonymous):

His book collection contains 24 textbooks. He can fit 18 books in one crate and 6 others in a second crate. Assuming the first crate is filled,how much space is left in the second crate?

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

what was the answer last time? to the size of crate?

OpenStudy (anonymous):

2520 in^3

OpenStudy (anonymous):

2520/18=140 each book is 140 in^3 So if he put 6 in second crate that means he took up 140*6=840 in^3 so 2520-840=1680 in^3 <- answer

OpenStudy (anonymous):

thanks:)))))